### v3

# Percolated stochastic block model via the EM algorithm and belief propagation with non-backtracking spectra

Abstract
Whereas Laplacian and modularity based spectral clustering is apt to dense graphs, recent results show that for sparse ones, the non-backtracking spectrum is the best candidate to find assortative clusters of nodes. Here belief propagation in the sparse stochastic block model is derived with arbitrary given model parameters that results in a non-linear system of equations; with linear approximation, the spectrum of the non-backtracking matrix is able to specify the number $k$ of clusters. Then the model parameters themselves can be estimated by the EM algorithm. Bond percolation in the assortative model is considered in the following two senses: the within- and between-cluster edge probabilities decrease with the number of nodes and edges coming into existence in this way are retained with probability $\beta$. As a consequence, the optimal $k$ is the number of the structural real eigenvalues (greater than $\sqrt{c}$, where $c$ is the average degree) of the non-backtracking matrix of the graph. Assuming, these eigenvalues $\mu_{1}>\dots>\mu_{k}$ are distinct, the multiple phase transitions obtained for $\beta$ are $\beta_{i}=\frac{c}{\mu_{i}^{2}}$; further, at $\beta_{i}$ the number of detectable clusters is $i$, for $i=1,\dots,k$. Inflation–deflation techniques are also discussed to classify the nodes themselves, which can be the base of the sparse spectral clustering.
Keywords: stochastic block model, belief propagation, EM algorithm, non-backtracking spectra, bond percolation.
Mathematics Subject Classification: 05C50, 05C80, 62H30
1 Introduction
In [9] we considered generalized quasirandom properties of expanding quasirandom graph sequences, which are deterministic counterparts of generalized random graphs, where the probabilities of edges coming into existence only depend on the cluster memberships of the endpoints, and are fixed during the expansion. If the number of the underlying clusters (number of nodes of the so-called model graph of [20]) is $k$, then a dense graph sequence is obtained in this way (the average degree of nodes is proportional to $n$, where $n$ is the number of nodes at a stage of the expansion). We proved that the adjacency matrix of such a random or quasirandom graph $G_{n}$ coming from the $k$-cluster model has $k$ structural eigenvalues (of order $n$), while the others are $o(n)$ in absolute value. Also, the normalized modularity matrix defined in [5] (its spectrum is in [-1,1]) has $k$ eigenvalues separated from 0, whereas the others are $o(1)$. Further, the clusters can be recovered by the weighted $k$-means algorithm applied to the $(k-1)$-dimensional representatives of the nodes obtained via the eigenvectors corresponding to the $k-1$ non-trivial structural eigenvalues (more precisely, the objective function of the $k$-means clustering is $o(1)$). In case of expanding regular graphs, the coordinates of any adjacency eigenvector (except the trivial one) are concentrated around certain values (according to a Gaussian distribution), keeping the orthogonality, see [2]. Also, in case of sparse graphs, the eigenvectors of the usual graph-based matrices are concentrated on the high degree nodes.
Unfortunately, the adjacency, Laplacian, or modularity matrix based methodology is not applicable in the sparse case, see [29]. To consider this situation, the edge probabilities are rescaled, with $n$ or $\log n$. These possibilities are discussed in [1]. In [23, 24], the authors consider scaling with $n$, and refine the block model threshold conjecture; they specify when information theoretically is possible or impossible to distinguish the clusters, depending on the model parameters. This so-called Kesten–Stigum threshold will be discussed in Section 2. Bickel and Chen [4] give a nonparametric statistical view of network models, and consider consistency in the strong or weak sense, when the estimated cluster assignment approaches the true one as $n\to\infty$. They prove that the likelihood modularity is always consistent, while the Newman–Girvan modularity is consistent only under stronger assumptions. They also consider different scalings. Other papers, instead of consistency, define different measures for the agreement of the true and the recovered clusterings (correlation, proportions of misclassifications, and information theoretical measures, e.g., the Kullbach–Leibler divergence). In the Laplacian or modularity based spectral clustering [6], $k$ is specified via spectral gaps, and the distance between the eigen-subspace spanned by the eigenvectors corresponding to the structural eigenvalues and that of the step-vectors is estimated by the Davis–Kahan theorem. As the square of this distance is the objective function of the k-means algorithm, this is at our disposal to find approximately the clusters in the dense case. This concept can be extended to non-symmetric matrices, like to the non-backtracking one, and due to theorems of [29] using Bauer–Fike type perturbation results, similar subspace perturbations are applicable to sparse graphs too.
In the sparse case we scale with $n$, so the average degree of the graph is of constant order. The so obtained sparse stochastic block model (called $SBM_{k}$) is assortative in the sense that the within-cluster probability of two nodes being connected is higher than the between-cluster one. Here bond percolation has two meanings: the first is that the within- and between-cluster edge probabilities decrease with $n$, and the second is that edges coming into existence in this way are retained with probability $\beta$. A real-life social network is an instance of such an expanding sequence, where two persons are connected randomly, with probability depending on their cluster memberships and this probability is the “smaller” the “larger” the network is; further, along these connections an information (e.g., an infection) is transmitted only with probability $\beta$ (which is related to the seriousness of the epidemic, akin to the recombination ratio).
The question is that for which values of the model parameters and $\beta$ the $k$ clusters can be distinguished with high probability (w.h.p.). If we knew $k$, the model parameters could be estimated with the EM algorithm, see [6]. The method of belief propagation (BP) is able to relate the model parameters to the non-backtracking spectrum of the graph, see [14, 17, 18]. However, the detectability thresholds have been given only in special cases (e.g., in the symmetric case of [14, 22], to be described in Section 2). In the present paper, BP is derived in the generic case; it provides a non-linear system of equations for the conditional probabilities of the nodes belonging to the clusters (given their neighbors), where the number of equations is $2mk$ ($m$ is the number of edges in the underlying graph). Though, there are numerical algorithms to solve this system of equations, BP is mainly important from theoretical point of view: via linearization and stability issues, our Theorem 1 establishes a close relation between the model parameters and the structural eigenvalues of the non-backtracking matrix $\bm{B}$ of the graph under mild assumptions. In this way, the number $k$ of clusters is approximately equal to the number of the structural eigenvalues of $\bm{B}$; but for given $k$, it is the EM algorithm that has the massive theoretical background to estimate the model parameters themselves.
We also experienced that, surprisingly, the leading $k$ (structural) eigenvalues of the expected adjacency matrix of a random graph coming from the sparse $SBM_{k}$ model are w.h.p. closer to the structural eigenvalues of $\bm{B}$ than to those of the adjacency matrix $\bm{A}$ of a randomly generated graph from this model, under some balancing conditions for the cluster sizes and average degrees of the clusters. This seems to contradict to the laws of large numbers, but in the sparse case (possessing nearly constant average degrees) it is supported by theoretical considerations (see, e.g., [29]).
Since the structural eigenvalues of $\bm{B}$ and the expected adjacency matrix are closely aligned, the so-called detectability threshold and $\beta$ also depend on these eigenvalues. However, in practical situations, only the non-backtracking spectrum is at our disposal and we do not know the model parameters. Based on this, we conclude the value of $k$ and derive multiple phase transitions in the general $SBM_{k}$ model, where, in addition, the edges of the graph coming from this model are retained with probability $\beta$, and we have the further percolated sparse stochastic block model $SBM_{k}^{\beta}$.
The rest of the paper is organized as follows. In Section 2, properties of the non-backtracking matrix are discussed together with basic notions and notation; further, the bond-percolation models are introduced. In Section 3, we perform BP theoretically which shows that $k$ is about the number of the structural eigenvalues of the non-backtracking matrix $\bm{B}$ that are larger than $\sqrt{c}$, where $c$ is the average degree of the random graph $G_{n}\in SBM_{k}$; it is Theorem 1 that supports this choice of $k$. In Section 5, we run the EM algorithm to estimate the parameters of the $SBM_{k}$ model that fits to our observed graph $G_{n}$. In the possession of the new estimates for the clusters and their average degrees, we adjust the number $k$ so that possible eigenvalues of $\bm{B}$ around $\sqrt{c}$ are added or deleted, and $k_{0}$ denotes the new number of clusters. In Section 4, we use inflation–deflation techniques for this purpose and apply a theorem of [29] to classify the majority of nodes by the k-means algorithm; this issue is summarized in Theorem 2 and a corollary and remark after it. Eventually, we find number $k$ of phase transitions for $\beta$ in the $SBM_{k}^{\beta}$ model, when $1,2,\dots,k$ clusters can be detected, see Section 6. Section 7 is devoted to conclusions and further perspectives.
2 Preliminaries
2.1 Non-backtracking matrix
The Hashimoto edge-incidence matrix, with other words, non-backtracking matrix $\bm{B}$ of a simple graph $G$ on $n$ nodes and $m$ edges is a $2m\times 2m$ non-symmetric matrix of 0-1 entries, see [17, 18, 21] in context of non-backtracking random walks, community detection, and centrality. The general entry of $\bm{B}$ is
$b_{ef}=\delta_{e\to f}\delta_{f\neq e^{-1}},$
for $e,f\in E^{\rightarrow}$, where $E^{\rightarrow}$ is the set of bidirected edges of $G$ (each existing edge is considered in both possible directions), and for $e=(e_{1},e_{2})$ the reversely directed edge is denoted $e^{-1}$, so $e^{-1}=(e_{2},e_{1})$; further, the $e\to f$ relation means that $e_{2}=f_{1}$ and $\delta$ is the (1-0) indicator of the event in its lower index. Therefore, $b_{ef}=1$ exactly when $e_{2}=f_{1}$ and $f_{2}\neq e_{1}$.
With other notation (used in Section 3) referring to the nodes, we have
$b_{l\rightarrow s,\,j\rightarrow i}=\delta_{sj}(1-\delta_{il}),$
where $\delta$ is now the Kronecker-delta. So $b_{l\rightarrow s,\,j\rightarrow i}=1$ if and only if for the quadruple in the lower indices, $l\rightarrow s=j\rightarrow i$ holds with $l\neq i$. Since it is a Frobenius-type matrix, its largest absolute value eigenvalue $\lambda(\bm{B})$ is positive real, and it can also have some other structural real eigenvalues (those are positive in assortative networks). Since the characteristic polynomial of $\bm{B}$ has real coefficients, its complex eigenvalues occur in conjugate pairs in the bulk of its spectrum. Note that the underlying simple graph is not directed, just we consider its edges as bidirected for the purpose of the message passing equations of Section 3.
Here we enumerate some results about the spectral properties of $\bm{B}$, see [16, 31] for more details and proofs. Note that the spectrum of $\bm{B}$ is not sensitive to high-degree nodes, because a walk starting at a node cannot return to it immediately. Also trees, disconnected from the graph or dangling off it, contribute zero eigenvalues to the non-backtracking spectrum, as a non-backtracking random walk when forced to a leaf, must stop there.
  • The frequently used Ihara’s formula (see [16, 18, 21]) implies that $\bm{B}$ has $m-n$ eigenvalues equal to 1 and $m-n$ eigenvalues equal to $-1$, whereas its further eigenvalues are those of the $2n\times 2n$ matrix
    $\bm{K}=\begin{pmatrix}\bm{O}&\bm{D}-\bm{I}_{n}\\ -\bm{I}_{n}&\bm{A}\end{pmatrix},$ (1)
    where $\bm{A}$ is the adjacency- and $\bm{D}$ is the diagonal degree-matrix of the graph (it contains the row-sums of $\bm{A}$, i.e., the node-degrees of $G$ in its main diagonal).
    Note that $\bm{K}$ always has at least one additional eigenvalue 1, the geometric multiplicity of which is equal to the number of the connected components of the underlying graph.
  • For the largest (positive real) eigenvalues of $\bm{A}$ and $\bm{B}$, the relation
    $\lambda(\bm{B})\leq\lambda(\bm{A})$ (2)
    holds (see [21]).
  • Though $\bm{B}$ is not a normal matrix, even not always diagonalizable (the algebraic and geometric multiplicity of some of its eigenvalues may not be the same), it exhibits some symmetry. Observe that
    $b^{*}_{ef}=b_{fe}=b_{e^{-1}\,f^{-1}}$
    for each directed pair of edges and for every entry $b^{*}$ of the transpose $\bm{B}^{*}$ of $\bm{B}$. This phenomenon can be described by involution and swapping.
    In the sequel, the vectors are column vectors and $*{}^{*}$ denotes the adjoint of a matrix or a vector (in case of real entries, it is the usual transposition).
  • Introduce the notation
    ${\breve{x}}_{e}:=x_{e^{-1}},\quad e\in E^{\rightarrow}$
    for relating the coordinates of the $2m$-dimensional vectors $\mathbf{x}$ and $\breve{\mathbf{x}}$ of $\mathbb{R}^{E^{\rightarrow}}$. Now $\mathbf{x}$ is partitioned into two $m$-dimensional vectors $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, where the coordinates of $\mathbf{x}_{1}$ correspond to $j\to i$ edges with $j<i$ and those of $\mathbf{x}_{2}$ correspond to their inverses. If $\mathbf{x}=(\mathbf{x}_{1}^{*},\mathbf{x}_{2}^{*})^{*}$, then ${\breve{\mathbf{x}}}=(\mathbf{x}_{2}^{*},\mathbf{x}_{1}^{*})^{*}$, i.e., $\mathbf{x}$ and $\breve{\mathbf{x}}$ can be obtained from each other by swapping the set of their first $m$ and second $m$ coordinates.
    Let $\bm{V}$ denote the following involution on $\mathbb{R}^{2m}$ ($\bm{V}=\bm{V}^{-1}$, $\bm{V}^{2}=\bm{I}$):
    $\bm{V}=\begin{pmatrix}\bm{O}&\bm{I}_{m}\\ \bm{I}_{m}&\bm{O}\end{pmatrix},$
    where the blocks are of size $m\times m$. With it, $\bm{V}\mathbf{x}={\breve{\mathbf{x}}}$ and $\bm{V}{\breve{\mathbf{x}}}=\mathbf{x}$.
    Proposition 1.
    Let us partition the non-backtracking matrix $\bm{B}$ of the connected simple graph $G$ on $n$ nodes and $m$ edges into the following four $m\times m$ quadrants:
    $\bm{B}=\begin{pmatrix}\bm{B}_{11}&\bm{B}_{12}\\ \bm{B}_{21}&\bm{B}_{22}\end{pmatrix},$
    where the first $m$ rows (columns) correspond to the $j\to i$ edges with $j<i$, and the next $m$ rows (columns) to their inverses. Then
[…]
    Further, $\bm{B}_{11}+\bm{B}_{12}+\bm{B}_{21}+\bm{B}_{22}$ is equal to the $m\times m$ adjacency matrix of the line-graph of $G$.
    Note that the nodes of the line-graph of $G$ are its edges, and two edges are adjacent if and only if they have a node in common. Let $\bm{N}$ denote the $n\times m$ node–edge (0-1) incidence matrix of the simple graph $G$ (except for trees, $n\leq m$). Then $\bm{N}\bm{N}^{*}$ and $\bm{N}^{*}\bm{N}$ are Gramians (covariance matrices), so they are positive semidefinite, have the same rank and the same set of positive eigenvalues. $\bm{N}\bm{N}^{*}=\bm{D}+\bm{A}$ is called signless Laplacian (whereas, the Laplacian matrix is $\bm{D}-\bm{A}$, see [6]). On the other hand, $\bm{N}^{*}\bm{N}-2\bm{I}_{m}$ is the adjacency matrix of the line-graph of $G$ (because the valence of each edge is 2). Consequently, its eigenvalues are those of the positive eigenvalues of $\bm{D}+\bm{A}$ decreased by 2, and the others are equal to -2, see [12] for more explanation. Furthermore, in [19] it is proved that if the line-graphs of two simple graphs, provided they both have node-degrees at least 4, are isomorphic, then they are isomorphic too. However, if the degree condition does not hold, it can happen that two not isomorphic graphs have isomorphic line-graphs. For example, if we have a triangle and a star on 4 vertices, then the adjacency matrix of their line-graphs is the same $3\times 3$ matrix (adjacency matrix of a complete graph, as any two edges of them have a node in common). The same happens if we have graphs with such dangling triangles or stars. Therefore, in case of sparse graphs, $\bm{B}$ carries more information for the graph than its line-graph; further, considering its edges in both directions makes it possible to perform non-backtracking random walks and message passing along them, see Section 3.
    Proof.
    (Proposition 1)
    Let us label the rows (and columns) of $\bm{B}$ with the edges, ordered lexicographically: the first $m$ rows correspond to the $j\to i$ edges $(1\leq j<i\leq n)$, whereas the second $m$ ones to the inverse edges $i\to j$, in the same order. Obviously, $\bm{B}_{12}$ is a symmetric matrix as its corresponding upper- and lower-diagonal entries, producing an entry 1, contain the matching edges in both directions; $\bm{B}_{21}$ is a symmetric matrix for the same reason. $\bm{B}_{11}$ is upper-, $\bm{B}_{22}$ is lower-triangular, and they are transposes of each other, because there is a one-to one correspondence between the upper-diagonal entries equal to 1 of $\bm{B}_{11}$ and the lower-diagonal entries equal to 1 of $\bm{B}_{22}$ such that one contains an $e\to f$ and the other an $f\to e$ relation. The other entries are equal to 0.
    The above argument shows that in the same position, at most only one entry can be equal to 1 in the four quadrants of $\bm{B}$, which means that the corresponding adjacency entry of the line-graph of $G$ is 1; the others are zeros. It is also obvious that the undirected $\{l,j\}$ and $\{j,i\}$ edges can be joined at $j$ in four possible ways: $i\to j\to l$, $l\rightarrow j\leftarrow i$, $i\leftarrow j\rightarrow l$, and $l\to j\to i$. Only one of them results in the true relation $\{l\to j\}\to\{j\to i\}$. This proves the second statement. The sum of the four quadrants of $\bm{B}$ is of course symmetric, because of the first statement. ∎
    As a consequence of Proposition 1,
    $(\bm{B}^{\ell})^{*}\bm{V}=\bm{V}\bm{B}^{\ell}$
    holds true for any natural number $\ell$. Applying this for $\ell=1$, the relation $\bm{B}^{*}=\bm{V}\bm{B}\bm{V}$ and
    $\bm{B}^{*}{\breve{\mathbf{x}}}=\breve{(\bm{B}\mathbf{x})}$
    is valid for any vector $\mathbf{x}\in\mathbb{R}^{2m}$. This implies the following: if $\mathbf{x}$ is a right eigenvector of $\bm{B}$ with a real eigenvalue $\mu$, then $\breve{\mathbf{x}}$ is a right eigenvector of $\bm{B}^{*}$ with the same eigenvalue. Consequently, if $\mathbf{x}$ is a right eigenvector of $\bm{B}$, then $\breve{\mathbf{x}}$ is a left eigenvector of it (and vice versa), with the same real eigenvalue.
  • But what is the relation between the eigenvectors corresponding to the leading (real) eigenvalues of $\bm{B}$ and $\bm{K}$? To ease the discussion, two auxiliary matrices, defined in [29] will be used: the $2m\times n$ end matrix $\bm{End}$ has entries $end_{ei}=1$ if $i$ is the end-node of the (directed) edge $e$ and 0, otherwise; the $2m\times n$ start matrix $\bm{Start}$ has entries $start_{ei}=1$ if $i$ is the start-node of the (directed) edge $e$ and 0, otherwise. Then for any vector $\mathbf{u}\in\mathbb{R}^{n}$ and for any edge $e=\{i\to j\}$ the following holds:
    $(\bm{End}\,\mathbf{u})(e)=u_{j}\quad\textrm{and}\quad(\bm{Start}\,\mathbf{u})(e)=u_{i}.$
    Consequently, $\bm{End}\,\mathbf{u}$ is the $2m$-dimensional inflated version of the $n$-dimensional vector $\mathbf{u}$, where the coordinate $u_{j}$ of $\mathbf{u}$ is repeated as many times, as many edges have end-node $j$; likewise, in the $2m$-dimensional inflated vector $\bm{Start}\,\mathbf{u}$, the coordinate $u_{i}$ of $\mathbf{u}$ is repeated as many times, as many edges have start-node $i$. As each edge is considered in both possible directions, this numbers are the node-degrees $d_{j}$ and $d_{i}$, respectively. Note that
    $\bm{End}^{*}\,\bm{End}=\bm{Start}^{*}\,\bm{Start}=\mathrm{diag}(d_{1},\dots,d_{n})=\bm{D}.$ (4)
    Akin to [18], for any vector $\mathbf{x}\in\mathbb{R}^{2m}$ define
    $x_{i}^{out}:=\sum_{j:\,j\sim i}x_{i\to j}\quad\textrm{and}\quad x_{i}^{in}:=\sum_{j:\,j\sim i}x_{j\to i},\quad i=1,\dots,n.$
    We put these coordinates into the $n$-dimensional (column) vectors $\mathbf{x}^{in}$ and $\mathbf{x}^{out}$. Trivially,
    $\mathbf{x}^{out}=\bm{Start}^{*}\mathbf{x},\quad\textrm{and}\quad\mathbf{x}^{in}=\bm{End}^{*}\mathbf{x},\quad i=1,\dots,n.$ (5)
    For them, we have the following relations (now more comfortably, we use $\bm{B}^{*}$ instead of $\bm{B}$):
    $(\bm{B}^{*}\mathbf{x})^{out}_{i}$ $=\sum_{e:\,e_{1}=i}(\bm{B}^{*}\mathbf{x})_{e}=\sum_{e:\,e_{1}=i}\sum_{f\to e,\,f\neq e^{-1}}x_{f}$ (6)
    $=\sum_{e:\,e_{1}=i}[\sum_{f\to e}x_{f}-x_{e^{-1}}]$
    $=\sum_{f:\,f_{2}=i}x_{f}\sum_{e:\,e_{1}=i}1-\sum_{e:\,e_{1}=i}x_{e^{-1}}$
    $=x_{i}^{in}d_{i}-\sum_{e:\,e^{-1}_{2}=i}x_{e^{-1}}=d_{i}x_{i}^{in}-x_{i}^{in}=(d_{i}-1)x_{i}^{in}.$
    Also,
    $(\bm{B}^{*}\mathbf{x})^{in}_{i}$ $=\sum_{e:\,e_{2}=i}(\bm{B}^{*}\mathbf{x})_{e}=\sum_{e:\,e_{2}=i}\sum_{f\to e,\,f\neq e^{-1}}x_{f}$ (7)
    $=\sum_{j=1}^{n}a_{ji}\sum_{f:\,f_{2}=j,\,f_{1}\neq i}x_{f}$
    $=\sum_{j=1}^{n}a_{ji}\sum_{f:\,f_{2}=j}x_{f}-\sum_{j=1}^{n}a_{ji}x_{i\to j}$
    $=\sum_{j=1}^{n}a_{ij}x^{in}_{j}-\sum_{j:\,j\sim i}x_{i\to j}=(\bm{A}\mathbf{x}^{in})_{i}-x^{out}_{i},$
    where we used that the (01)(0--1)( 0 - - 1 ) adjacency matrix $\bm{A}$ of the graph is symmetric with entries $a_{ij}=a_{ji}=\delta_{i\sim j}$.
    Summarizing, if $\mathbf{x}$ is a right eigenvector of $\bm{B}^{*}$ with (real) eigenvalue $\mu$, then
    $\mu\begin{pmatrix}\mathbf{x}^{out}\\ \mathbf{x}^{in}\end{pmatrix}=\begin{pmatrix}(\bm{B}^{*}\mathbf{x})^{out}\\ (\bm{B}^{*}\mathbf{x})^{in}\end{pmatrix}=\begin{pmatrix}\bm{O}&\bm{D}-\bm{I}_{n}\\ \bm{-}\bm{I}_{n}&\bm{A}\end{pmatrix}\begin{pmatrix}\mathbf{x}^{out}\\ \mathbf{x}^{in}\end{pmatrix}$ (8)
    and
    $\begin{pmatrix}(\bm{B}^{*}\mathbf{x})^{out}\\ (\bm{B}^{*}\mathbf{x})^{in}\end{pmatrix}=\begin{pmatrix}(\bm{D}-\bm{I}_{n})\mathbf{x}^{in}\\ \bm{A}\mathbf{x}^{in}-\mathbf{x}^{out}\end{pmatrix}=\bm{K}\begin{pmatrix}\mathbf{x}^{out}\\ \mathbf{x}^{in}\end{pmatrix}.$
    In particular, if $\mathbf{x}$ is a right eigenvector of $\bm{B}^{*}$ with a real eigenvalue $\mu\neq 0$, then the $2n$-dimensional vector comprised of parts $\mathbf{x}^{out}$ and $\mathbf{x}^{in}$ is a right eigenvector of $\bm{K}$ with the same eigenvalue. Indeed,
    $\mu\begin{pmatrix}\mathbf{x}^{out}\\ \mathbf{x}^{in}\end{pmatrix}=\begin{pmatrix}(\mu\mathbf{x})^{out}\\ (\mu\mathbf{x})^{in}\end{pmatrix}=\begin{pmatrix}(\bm{B}^{*}\mathbf{x})^{out}\\ (\bm{B}^{*}\mathbf{x})^{in}\end{pmatrix}=\bm{K}\begin{pmatrix}\mathbf{x}^{out}\\ \mathbf{x}^{in}\end{pmatrix}.$
    According to the previous remarks, the vector $\mathbf{x}$ is a left eigenvector, and $\breve{\mathbf{x}}$ is a right eigenvector of $\bm{B}$ with the same eigenvalue. To both of them the two segments, $\mathbf{x}^{out}$ and $\mathbf{x}^{in}$ of the right eigenvector of $\bm{K}$ are responsible. In view of the relation $\mathbf{x}^{out}=\frac{1}{\mu}(\bm{D}-\bm{I}_{n})\mathbf{x}^{in}$, obtained by (8), it suffices to consider only $\mathbf{x}^{in}\in\mathbb{R}^{n}$ for further clustering purposes, see Section 4.
2.2 Random graphs obtained by bond-percolation
In the simple graph scenario, in [26], it is proved that the so-called message passing system of equations (equivalently, BP) has a non-trivial solution if and only if $\beta>\frac{1}{\lambda(\bm{B})}$, where $\lambda(\bm{B})$ is the largest eigenvalue of the non-backtracking matrix $\bm{B}$ of a given sparse social network. So $\frac{1}{\lambda(\bm{B})}$ is the bond-percolation threshold for the giant component to appear in this simple case; which in view of (2) is greater than $\frac{1}{\lambda(\bm{A})}$ of [10] in the dense case.
Now our purpose is to characterize the non-backtracking spectrum of a random graph $G_{n}$ on $n$ nodes coming from the $SBM_{k}$ model, and via perturbations, try to use these properties to find out the number of clusters and the clusters themselves. The parameters of the sparse $SBM_{k}$ model are as follows.
The $k\times k$ probability matrix $\bm{P}$ of the percolated sparse random graph $G_{n}\in SBM_{k}$ has entries
$p_{ab}=\frac{c_{ab}}{n},$
where the $k\times k$ symmetric affinity matrix $\bm{C}=(c_{ab})$ stays constant as $n\to\infty$. An edge between $i<j$ comes into existence, independently of the others, with probability $p_{ab}$ if $i\in V_{a}$ and $j\in V_{b}$, where $(V_{1},\dots,V_{k})$ is a partition of the node-set $V$ into $k$ disjoint clusters. This will produce the upper-diagonal part of the $n\times n$ random adjacency matrix $\bm{A}$, and $a_{ji}:=a_{ij}$. It can be extended to the $i=j$ case when self-loops are allowed, or else, the diagonal entries of the adjacency matrix are zeros.
Let $\bar{\bm{A}}$ denote the $n\times n$ inflated matrix of the $k\times k$ matrix $\bm{P}$: ${\bar{a}}_{ij}=p_{ab}$ if $i\in V_{a}$ and $b\in V_{b}$. When loops are allowed, then $\mathbb{E}(a_{ij})={\bar{a}}_{ij}$ for all $1\leq i,j\leq n$. In the loopless case, the expected adjacency matrix $\mathbb{E}\bm{A}$ differs from $\bar{\bm{A}}$ with respect to the the main diagonal, but the diagonal entries are negligible as will be shown in Section 4.
Specifically, sometimes $c_{ab}=c_{in}$ is the within-cluster $(a=b)$ and $c_{ab}=c_{out}$ is the between-cluster $(a\neq b)$ affinity. In [25], the network is called assortative if $c_{in}>c_{out}$, and disassortative if $c_{in}<c_{out}$. Of course, remarkable difference is needed between the two, to recognize the clusters. The cluster sizes are $n_{1},\dots,n_{k}$ ($\sum_{i=1}^{k}n_{i}=n$), so the $k\times k$ diagonal matrix $\bm{R}:=\mathrm{diag}(r_{1},\dots,r_{k})$, where $r_{a}=\frac{n_{a}}{n}$ is the relative size of cluster $a$ $(a=1,\dots,k)$, is also a model parameter ($\sum_{a=1}^{k}r_{a}=1$). The model $SBM_{k}$ is called symmetric if $r_{1}=\dots=r_{k}=\frac{1}{k}$ and all diagonal entries of the affinity matrix are equal to $c_{in}$, whereas the off-diagonal ones to $c_{out}$.
The average degree of a real world graph on $m$ edges and $n$ nodes is $\frac{2m}{n}$. The expected average degree of the random graph $G_{n}$ generated from the $SBM_{k}$ model is
$c=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}{\bar{a}}_{ij}=\frac{1}{n}\sum_{a=1}^{k}\sum_{b=1}^{k}n_{a}n_{b}p_{ab}=\frac{1}{n^{2}}\sum_{a=1}^{k}\sum_{b=1}^{k}n_{a}n_{b}c_{ab}=\sum_{a=1}^{k}r_{a}c_{a},$ (9)
where $c_{a}=\sum_{b=1}^{k}r_{b}c_{ab}$ is the average degree of cluster $a$. It is valid only if self-loops are allowed. Otherwise, $c_{a}$ and $c$ should be decreased with a term of order $\frac{1}{n}$, but it will not make too much difference in the subsequent calculations.
In [11], the case when $c_{a}=c$ for all $a$ is considered. In this case $\frac{1}{c}{\bar{\bm{A}}}$ is a stochastic matrix, and so, the spectral radius of $\bar{\bm{A}}$ is $c$.
The symmetric case is a further special case of this, when
$c=\frac{c_{in}+(k-1)c_{out}}{k}.$
In this case, the separation of the clusters only depends on the $c_{in}$, $c_{out}$ relation. If $c_{in}$ is “close” to $c_{out}$, then the groups cannot be distinguished. The detectability threshold (Kesten–Stigum threshold) in the symmetric case is
$|c_{in}-c_{out}|>k\sqrt{c},$ (10)
see [11, 14, 22].
In the symmetric $k=2$ case, the authors of [23] prove that under condition (10), their algorithm gives a clustering that has a correlation with the true one separated from zero; furthermore, this correlation tends to 1 as the ratio of the left and right hand sides of (10) increases. They also write that the $k\geq 2$ and $c_{1}=\dots=c_{k}$ case is the hardest one, as otherwise the clusters could be distinguished by sorting the node-degrees.
Note that Theorem 4 of [11] allows a small fluctuation for $n_{a}$’s, and so, for $r_{a}$’s and $c_{a}$’s too. We make use of this when compare the structural eigenvalues of the non-backtracking matrix to the eigenvalues of the so-called transition matrix assessed during the BP process. Observe that these assumptions are natural in social networks: the relative sizes of communities are balanced (tend to a limit as $n\to\infty$); further, the average degrees of nodes of the clusters do not differ much as $n\to\infty$, which means that members of different communities have on average closely the same number of bonds to members of their own or other communities.
Stephan and Massoulié [29] investigate more general random graph models, also in the edge-weighted case. In the special unweighted situation we shall use the following (informal) statement of them for the $SBM_{k}$ model, where the matrix $\bar{\bm{A}}$ is not only the diagonal-corrected expected adjacency matrix, but it is approximately the variance matrix of the random adjacency matrix $\bm{A}$.
Proposition 2 (based on Theorem 1 of [29]).
Assume that $k=\mathrm{rank}\,{\bar{\bm{A}}}=n^{o(1)}$, the graph is sparse enough, and the eigenvectors corresponding to the non-zero eigenvalues of the matrix $\bar{\bm{A}}$ are sufficiently delocalized. Let $k_{0}$ denote the number of eigenvalues of $\bar{\bm{A}}$ whose absolute value is larger than $\sqrt{\rho}$, where $\rho$ is the spectral radius of $\bar{\bm{A}}$: these are $\nu_{1}\geq\dots\geq\nu_{k_{0}}$ with corresponding eigenvectors $\mathbf{u}_{1},\dots,\mathbf{u}_{k_{0}}$ (they form an orthonormal system as $\bar{\bm{A}}$ is a real symmetric matrix). Then for $i\leq k_{0}\leq k$, the $i$th largest eigenvalue $\mu_{i}$ of $\bm{B}$ is asymptotically (as $n\to\infty$) equals to $\nu_{i}$ and all the other eigenvalues of $\bm{B}$ are constrained to the circle (in the complex plane) of center 00 and radius $\sqrt{\rho}$. Further, if $i\leq k_{0}$ is such that $\nu_{i}$ is a sufficiently isolated eigenvalue of $\bar{\bm{A}}$, then the standardized eigenvector of $\bm{B}$ corresponding to $\mu_{i}$ has inner product close to 1 with the standardized inflated version of $\mathbf{u}_{i}$, namely, with $\frac{\bm{End}\,\mathbf{u}_{i}}{\|\bm{End}\,\mathbf{u}_{i}\|}$.
Note that if our random graph comes from the $SBM_{k}$ model with $k\times k$ parameter matrices $\bm{R}$ and $\bm{C}$, then $\bar{\bm{A}}$ has rank at most $k$, and its nonzero eigenvalues are identical to the real eigenvalues of the matrix $\bm{R}\bm{C}$ (same as those of $\bm{R}^{\frac{1}{2}}\bm{C}\bm{R}^{\frac{1}{2}}$), and the eigenvectors of it are inflated versions of those of $\bm{R}^{\frac{1}{2}}\bm{C}\bm{R}^{\frac{1}{2}}$; consequently, they are step-vectors on $k$ different steps and so, the k-means algorithm is applicable to the $k_{0}$-dimensional representatives of the nodes constructed with the $k_{0}$ normalized eigenvectors of $\bm{B}$ corresponding to its leading eigenvalues $\mu_{1},\dots,\mu_{k_{0}}$. The point is that we do not need the model parameters and the spectral decomposition of $\bar{\bm{A}}$ itself for the k-means clustering. This issue together with inflation–deflation techniques will be discussed in details in Section 4. Also note that in the $SBM_{k}$ model, when $c_{a}=c$ $(a=1,\dots,k)$, then $\rho=c$ and $k_{0}=k$.
3 Belief propagation (BP) in the general SBM
To derive the percolation threshold analytically, we use the BP or cavity method of [22] in the generic case. In [22] it was developed for the symmetric case. Similarly to the BP, called message passing in [26], let us introduce the following notation: for $a=1,\dots,k$,
$\psi_{i}^{a}\propto\mathbb{P}(i\textrm{ is in the cluster }a),$
given the observed graph on $n$ nodes, coming from the $SBM_{k}$ model, where $\psi_{i}^{a}$ for $a=1,\dots,k$ defines the marginal membership (state) distribution of node $i$.
We assume that our neighbors are independent of each other, when conditioned on our own state. Equivalently, we assume that our neighbors are correlated only through us. This can be modeled by having each node $j$ send a message $\psi_{j\to i}$ to $i$, which is an estimate of $j$’s marginal (in other words, state or membership) if $i$ were not there (more precisely, if we did not know whether or not there is an edge between $i$ and $j$). Therefore, the conditional probability
$\psi_{j\to i}^{a}:=\mathbb{P}(j\textrm{ is in cluster }a\textrm{ when }i\textrm{ is not present})$
(or when we do not know the membership of $i$) can be computed through the neighbors of $j$ that are different from $i$ (in the realization of the random graph coming from the $SBM_{k}$ model) as follows:
$\psi_{j\to i}^{a}=C_{a}^{ij}r_{a}\prod_{l\sim j,\,l\neq i}\sum_{b=1}^{k}\psi_{l\to j}^{b}\,p_{ab},\quad a=1,\dots,k,$ (11)
where $C_{a}^{ij}$ is a normalizing factor. Here $\psi_{l\to j}^{b}$ is the conditional probability that node $l$ (which is not $i$ and to which $j$ is connected) belongs to cluster $b$ even in the absence of $j$, and the summation is an application of the rule of completely disjoint events with probabilities $p_{ab}$ for $b=1,\dots,k$.
This so-called BP or message-passing equation is a system of $2mk$ non-linear equations with the same number of unknowns. It can be solved by initializing messages randomly, then repeatedly updating them via (11). This procedure usually converges quickly and the resulting fixed point gives a good estimate of the marginals:
$\psi_{i}^{a}\propto r_{a}\prod_{j\sim i}\sum_{b=1}^{k}\psi_{j\to i}^{b}\,p_{ab},$
where the constant of proportionality is chosen according to $\sum_{a=1}^{k}\psi_{i}^{a}=1$. However, the system of equations contains the model parameters, so it can be solved only if the model parameters are known.
In [22], the symmetric case is treated, when BP has a trivial fixed point $\psi_{j\to i}^{a}=\frac{1}{k}$, for $a=1,\dots,k$. If it gets stuck there, then BP does no better than chance. It happens when this trivial fixed point of this discrete dynamical system is asymptotically stable.
In the generic case, we want to have an unstable fixed point via linearization:
$\psi_{j\to i}^{a}:=r_{a}+\varepsilon_{j\to i}^{a}.$
If we substitute this in (11) and expand to first order in $\bm{\varepsilon}$ (vector of $2mk$ coordinates $\varepsilon_{j\to i}^{a}$’s), the updated system of equations is
$\varepsilon_{j\to i}^{a}$ $=\psi_{j\to i}^{a}-r_{a}=r_{a}\left\{C_{a}^{ij}\prod_{l\sim j,\,l\neq i}\left[\sum_{b=1}^{k}\psi_{l\to j}^{b}\,p_{ab}\right]-1\right\}$ (12)
$=r_{a}\left\{C_{a}^{ij}\prod_{l\sim j,\,l\neq i}\left[\sum_{b=1}^{k}(r_{b}+\varepsilon_{l\to j}^{b})\,p_{ab}\right]-1\right\}$
$=r_{a}\left\{C_{a}^{ij}\prod_{l\sim j,\,l\neq i}\left[\sum_{b=1}^{k}r_{b}p_{ab}+\sum_{b=1}^{k}\varepsilon_{l\to j}^{b}\,p_{ab}\right]-1\right\}$
$=r_{a}\left\{C_{a}^{ij}\prod_{l\sim j,\,l\neq i}\left[\frac{c_{a}}{n}+\sum_{b=1}^{k}\varepsilon_{l\to j}^{b}\,\frac{c_{ab}}{n}\right]-1\right\}$
$=r_{a}\left\{C_{a}^{ij}(\frac{1}{n})^{s_{j}-1}\left[\sum_{b=1}^{k}\sum_{l\sim j,\,l\neq i}\varepsilon_{l\to j}^{b}\,c_{ab}\,c_{a}^{s_{j}-2}+c_{a}^{s_{j}-1}\right]-1\right\}+O(\bm{\varepsilon}^{2}),$
where $s_{j}$ denotes the number of neighbors of $j$ and $s_{j}-1$ is the number of its neighbors that are different from $i$ (this number is frequently 0 or 1, as we have a sparse graph). If $s_{j}<2$ happens, then the corresponding entry of the non-backtracking matrix is 0. To specify the normalizing factor $C_{a}^{ij}$, we substitute zeros for $\varepsilon$’s that provide the trivial solution. This approximately yields
$C_{a}^{ij}\left(\frac{1}{n}\right)^{s_{j}-1}\,c_{a}^{s_{j}-1}-1=0,$
so
$C_{a}^{ij}=\left(\frac{n}{c_{a}}\right)^{s_{j}-1}.$
Substituting this into the last equation of (12), we get
$\varepsilon_{j\to i}^{a}$ $=r_{a}\left\{\left(\frac{n}{c_{a}}\right)^{s_{j}-1}\left(\frac{1}{n}\right)^{s_{j}-1}c_{a}^{s_{j}-2}\left[\sum_{b=1}^{k}\sum_{l\sim j,l\neq i}\varepsilon_{l\to j}^{b}\,c_{ab}\,+c_{a}\right]-1\right\}+O(\bm{\varepsilon}^{2})$
$=r_{a}\left\{\frac{1}{c_{a}}\left[\sum_{b=1}^{k}\sum_{l\sim j,l\neq i}\varepsilon_{l\to j}^{b}\,c_{ab}\,+c_{a}\right]-1\right\}+O(\bm{\varepsilon}^{2}).$
The linear approximation of the above system of equations is
$\bm{\varepsilon}=\bm{M}\bm{\varepsilon},$ (13)
where $\bm{\varepsilon}$ is a $2mk$-dimensional vector and the $(j\to i,\,a)\,,\,(l\to s,\,b)$ entry of $\bm{M}$ is $\partial\varepsilon_{j\to i}^{a}/\partial\varepsilon_{l\to s}^{b}$. It is not zero only for $j\to i$, $l\to s$ pairs with $s=j$ and $l\neq i$, i.e., when the corresponding entry of $\bm{B}$ is 1. If it is not zero, then the partial derivative of $\varepsilon_{j\to i}^{a}$ with respect to $\varepsilon_{l\to j}^{b}$ is $\frac{r_{a}}{c_{a}}c_{ab}$ (the second order terms are disregarded). Therefore, it can easily be seen that $\bm{M}$ is a Kronecker-product:
$\bm{M}=\bm{B}\otimes\bm{T},$
i.e., $\bm{T}$ is substituted for each entry 1 of the non-backtracking matrix $\bm{B}$, where $\bm{T}=\bm{G}\bm{R}\bm{C}$ is the transmission matrix with $\bm{G}=\mathrm{diag}(\frac{1}{c_{1}},\dots,\frac{1}{c_{k}})$. Note that $\bm{T}$ is a stochastic matrix, so its largest eigenvalue is 1, and the others are less than 1 and positive in the assortative case.
For the solution of (13), the fixed point of the linear dynamical system
$\bm{\varepsilon}^{(t+1)}=\bm{M}\bm{\varepsilon}^{(t)},\quad t=1,2,\dots$
is looked for. We can find a fixed point other than the trivial $\mathbf{0}$ if $\mathbf{0}$ is an unstable solution, for which a sufficient condition is that the spectral radius of the matrix $\bm{M}$ is greater than 1. In this way, we have proved the following.
Theorem 1.
Assume that the graph $G$ comes from the $SBM_{k}$ model with positive integer $k$ and $k\times k$ parameter matrices $\bm{R}=\mathrm{diag}(r_{1},\dots,r_{k})$ (cluster proportions) and $\bm{C}$ (symmetric affinity matrix). The approximating linear dynamical system behind the BP system (12) is $\bm{\varepsilon}^{(t+1)}=\bm{M}\bm{\varepsilon}^{(t)}$, where $\bm{\varepsilon}$ is a $2mk$-dimensional vector and the $2mk\times 2mk$ matrix of the system is $\bm{M}=\bm{B}\otimes\bm{T}$. Here $\bm{B}$ is the non-backtracking matrix of the graph $G$ and $\bm{T}=\bm{G}\bm{R}\bm{C}$ is the transmission matrix with $\bm{G}=\mathrm{diag}(\frac{1}{c_{1}},\dots,\frac{1}{c_{k}})$, where $c_{a}=\sum_{b=1}^{k}r_{b}c_{ab}$ is the average degree of cluster $a$, for $a=1,\dots,k$. The trivial fixed point $\mathbf{0}$ is an unstable fixed point of the above system if there is at least one eigenvalue of the matrix $\bm{B}\otimes\bm{T}$ (product of eigenvalues of $\bm{B}$ and $\bm{T}$) that is greater than 1 (in absolute value).
Note that in [22] only the symmetric case is treated. In [11] the special case $c_{1}=\dots=c_{k}=c$ is considered when the matrix $\bm{T}$ becomes $\bm{T}=\frac{1}{c}\bm{R}\bm{C}$; then the leading eigenvalues of $\bm{B}$ and $\bm{R}\bm{C}$ are w.h.p. “close” to each other. Also, the largest eigenvalue of $\bm{R}\bm{C}$ is $c$, which is trivially the case if $k=1$ and we have the Erdős–Rényi random graph. More precisely, the authors of [11] allow “small” fluctuations of the cluster membership proportions that causes the same order of fluctuations in the average degrees of the clusters. For the membership proportion of cluster $a$, denoted by $r_{a}^{(n)}$, the relation $\lim_{n\to\infty}r_{a}^{(n)}=r_{a}$ is assumed with $\sum_{a=1}^{k}r_{a}=1$. Specifically, in [11], the assumption
$\max_{a\in\{1,\dots,k\}}|r_{a}^{(n)}-r_{a}|=O(n^{-\gamma})$ (14)
is made with some $\gamma\in(0,1]$, where $n$ denotes the number of nodes. This assumption implies that in the $c_{1}=\dots=c_{k}=c$ case, $\max_{a\in\{1,\dots,k\}}|c_{a}^{(n)}-c|=O(n^{-\gamma})$.
Theorem 4 of [11] states that if
$\max_{a}c_{a}^{(n)}=c+O(n^{-\gamma}),$ (15)
with some $\gamma\in(0,1]$, and the relative proportions of the clusters converge, then w.h.p.
$\mu_{i}=\nu_{i}+o(1)\quad(i=1,\dots k_{0})\quad\textrm{and}\quad\mu_{i}<\sqrt{c}+o(1)\quad(i>k_{0}),$
where $\mu_{i}$’s and $\nu_{i}$’s $(i=1,\dots,k_{0})$ are the structural eigenvalues of $\bm{B}$ and $\bm{R}\bm{C}$, respectively, whereas $k_{0}\leq k$ is the positive integer for which $\nu_{i}^{2}\geq\nu_{1}$ $(i=1,\dots k_{0})$ and $\nu_{k_{0}+1}^{2}<\nu_{1}$ holds. Therefore, in the $SBM_{1}$ (Erdős–Rényi) model, $\mu_{1}=c+o(1)$ and $\mu_{2}\leq\sqrt{c}+o(1)$.
However, in the generic case too, simulations show that sometimes the leading eigenvalues of $\bm{B}$ and $\bm{R}\bm{C}$ are closer to each other than those of $\bm{B}$ and $c\bm{T}$, but always closer than those of $\bm{A}$ and any of them. The following considerations can be made. If (15) holds, then
$\min_{a}\frac{c}{c_{a}^{(n)}}=\frac{c}{\max_{a}c_{a}^{(n)}}=1+O(n^{-\gamma}).$ (16)
By the rules of the Kronecker-products,
$\bm{B}\otimes c\bm{T}=(\bm{I}_{2m}\bm{B})\otimes(c\bm{G}\bm{R}\bm{C})=(\bm{I}_{2m}\otimes c\bm{G})(\bm{B}\otimes\bm{R}\bm{C}).$
But
$\bm{I}_{2m}\otimes c\bm{G}=\mathrm{diag}(\frac{c}{c_{1}},\dots.\frac{c}{c_{k}},\dots,\frac{c}{c_{1}},\dots.\frac{c}{c_{k}})$
repeated $2m$ times. By minimax theorems, the eigenvalues of $(\bm{I}_{2m}\otimes c\bm{G})(\bm{B}\otimes\bm{R}\bm{C})$ are within factors $u$ and $v$ of those of $\bm{B}\otimes\bm{R}\bm{C}$, where
$u=\min_{a}\frac{c}{c_{a}^{(n)}}\quad\textrm{and}\quad v=\max_{a}\frac{c}{c_{a}^{(n)}}.$
Here $u$ and $v$ depend on $n$, but we do not denote this dependence. Also, for fixed “large” $n$, the eigenvalues of $c\bm{G}\bm{R}\bm{C}$ are within a factor $u$ and $v$ of those of $\bm{R}\bm{C}$.
Then the condition $\lambda(\bm{B})\,\lambda(\bm{T})>1$, i.e.,
$\lambda(\bm{B})\,\lambda(c\bm{T})=\lambda(\bm{B})\,\lambda(c\bm{G}\bm{R}\bm{C})>c$
holds if
$\lambda(\bm{B})\,(u\lambda(\bm{B}))>c.$
This implies that
$\lambda(\bm{B})\geq\frac{\sqrt{c}}{\sqrt{u}}\quad\textrm{and}\quad\lambda(c\bm{G}\bm{R}\bm{C})\geq u\lambda(\bm{B})\geq\sqrt{u}\sqrt{c}$
should hold, where $u\leq 1$.
Consequently, real eigenvalues of $\bm{B}$ that are larger than $\sqrt{c}$ should be taken into account. However, $\bm{R}$ and $\bm{C}$ contain the unknown model parameters, $\bm{G}$ depends on them, and there is only $\bm{B}$ at our disposal from a sample. But the model itself is adapted to the data, i.e., the underlying graph. In [26], the author makes the following crucial observation: “The calculation, we have described is based on a fit of the stochastic block model to a network that was itself generated from the same model.”
So we should consider the real eigenvalues of $\bm{B}$ that are greater than $\sqrt{c}$. This number will be denoted by $k_{0}$. Also, the non-zero eigenvalues of the expected adjacency matrix $\bar{\bm{A}}$ of the next section are the same as those of $\bm{R}\bm{C}$, so they are in the neighborhood of the leading eigenvalues of $\bm{B}$ within a factor between $u$ and $v$. On the contrary, their closeness to $\bm{A}$ is up to an additive constant (less than $\sqrt{c}$) as will be shown in Section 4.
Note that when the average degrees $c_{a}$’s of the clusters differ significantly, then the leading eigenvalues of $\bm{B}$ are not closely aligned with those of $\bm{R}\bm{C}$. However, this case is relatively simple, as then the clusters could be formed by sorting the node degrees.
4 Inflation–deflation
Here we use the inflation–deflation technique of [6].
With the notation of the $SBM_{k}$ model: $\bm{C}=(c_{ab})$ is the $k\times k$ symmetric affinity matrix and $\bm{P}=\frac{1}{n}\bm{C}$ is the corresponding probability matrix of the random graph $G_{n}\in SBM_{k}$. If self-loops are allowed, then $\bar{\bm{A}}$ is the expected adjacency matrix of $G_{n}$ and it contains $p_{ab}$’s in the $(a,b)$-th block. Therefore, it is the inflated matrix of $\bm{P}$ with blow-up sizes $n_{1},\dots,n_{k}$ $(\sum_{a=1}^{k}n_{a}=n)$. The blow-up ratios are $r_{a}=\frac{n_{a}}{n}$ for $a=1,\dots,k$. The expected average degree of $G_{n}$ is
$c=\frac{1}{n}\sum_{i=1}^{n}\sum_{i=1}^{n}{\bar{a}}_{ij}=\frac{1}{n}\sum_{a=1}^{k}\sum_{b=1}^{k}n_{a}n_{b}c_{ab}=\sum_{a=1}^{k}r_{a}c_{a},$
as in (9). Note that $c$ depends only on the model parameters and it is of constant order. If the diagonal entries of $\bm{A}$ are zeros, then $\bar{\bm{A}}$ differs from the expected adjacency matrix by $\mathrm{diag}(p_{11},\dots,p_{11},\dots,p_{kk},\dots,p_{kk})$ with $n_{a}$ entries equal to each other in the $a$-th block. However, by the Weyl’s perturbation theorem, the $k$ leading eigenvalues of $\bar{\bm{A}}$ and $\mathbb{E}(\bm{A})$ differ only with a term of order $\frac{1}{n}$.
Proposition 3.
The matrix ${\bar{\bm{A}}}$ has rank $k$ and its non-zero eigenvalues ($\nu$’s) with unit norm eigenvectors ($\mathbf{u}$’s) satisfy
${\bar{\bm{A}}}\mathbf{u}=\nu\mathbf{u},$ (17)
where $\mathbf{u}$ is the inflated vector of ${\tilde{\mathbf{u}}}=(u(1),\dots,u(k))^{*}$ with block-sizes $n_{1},\dots,n_{k}$. With the notation $\bm{R}=\frac{1}{n}\mathrm{diag}(n_{1},\dots,n_{k})=\mathrm{diag}(r_{1},\dots,r_{k})$, the deflated equation of (17) is equivalent to
$\bm{R}^{\frac{1}{2}}\bm{C}\bm{R}^{\frac{1}{2}}\mathbf{v}=\nu\mathbf{v},$ (18)
where $\mathbf{v}=\sqrt{n}\bm{R}^{\frac{1}{2}}{\tilde{\mathbf{u}}}$. Further, if $\mathbf{u}_{1},\dots\mathbf{u}_{k}$ is the set of orthonormal eigenvectors of $\bar{\bm{A}}$, then $\mathbf{v}_{i}=\sqrt{n}\bm{R}^{\frac{1}{2}}{\tilde{\mathbf{u}}}_{i}$ $(i=1,\dots,k)$ is the set of orthonormal eigenvectors of $\bm{R}^{\frac{1}{2}}\bm{C}\bm{R}^{\frac{1}{2}}$. Also, $\bm{R}^{\frac{1}{2}}\mathbf{v}_{i}=\sqrt{n}\bm{R}{\tilde{\mathbf{u}}}_{i}$ is a right eigenvector of $\bm{R}\bm{C}$ and $\bm{R}^{-\frac{1}{2}}\mathbf{v}_{i}=\sqrt{n}{\tilde{\mathbf{u}}}_{i}$ is a left eigenvector of $\bm{R}\bm{C}$ with eigenvalue $\nu_{i}$, for $i=1,\dots,k$.
Proof.
The matrix ${\bar{\bm{A}}}$ has $k$ different rows (columns), so it is of rank $k$ and its non-zero eigenvalues ($\nu$’s) with unit norm eigenvectors ($\mathbf{u}$’s) trivially satisfy (17), where $\mathbf{u}$ is the inflated vector of ${\tilde{\mathbf{u}}}=(u(1),\dots,u(k))^{*}$ with block-sizes $n_{1},\dots,n_{k}$. The deflated equation of (17) is equivalent to
$\bm{R}^{\frac{1}{2}}\bm{P}\bm{R}^{\frac{1}{2}}\mathbf{v}=\frac{\nu}{n}\mathbf{v},$
that in turn is equivalent to (18), where the eigenvector $\mathbf{v}=\sqrt{n}\bm{R}^{\frac{1}{2}}{\tilde{\mathbf{u}}}$ also has unit norm.
If $\mathbf{u}_{1},\dots\mathbf{u}_{k}$ is the set of orthonormal eigenvectors of $\bar{\bm{A}}$, then $\mathbf{v}_{i}=\sqrt{n}\bm{R}^{\frac{1}{2}}{\tilde{\mathbf{u}}}_{i}$ $(i=1,\dots,k)$ is the set of orthonormal eigenvectors of $\bm{R}^{\frac{1}{2}}\bm{C}\bm{R}^{\frac{1}{2}}$. Also,
$\bm{R}\bm{C}(\bm{R}^{\frac{1}{2}}\mathbf{v}_{i})=\bm{R}^{\frac{1}{2}}(\bm{R}^{\frac{1}{2}}\bm{C}\bm{R}^{\frac{1}{2}})\mathbf{v}_{i}=\bm{R}^{\frac{1}{2}}(\bm{R}^{\frac{1}{2}}\bm{C}\bm{R}^{\frac{1}{2}})\mathbf{v}_{i}=\nu_{i}\bm{R}^{\frac{1}{2}}\mathbf{v}_{i},$
so $\bm{R}^{\frac{1}{2}}\mathbf{v}_{i}=\sqrt{n}\bm{R}{\tilde{\mathbf{u}}}_{i}$ is a right eigenvector of $\bm{R}\bm{C}$ for $i=1,\dots,k$. Further,
$(\bm{R}^{-\frac{1}{2}}\mathbf{v}_{i})^{*}\bm{R}\bm{C}=\mathbf{v}_{i}^{*}(\bm{R}^{\frac{1}{2}}\bm{C}\bm{R}^{\frac{1}{2}})\bm{R}^{-\frac{1}{2}}=\nu_{i}\mathbf{v}_{i}^{*}\bm{R}^{-\frac{1}{2}}=\nu_{i}(\bm{R}^{-\frac{1}{2}}\mathbf{v}_{i})^{*},$
so $\bm{R}^{-\frac{1}{2}}\mathbf{v}_{i}=\sqrt{n}{\tilde{\mathbf{u}}}_{i}$ is a left eigenvector of $\bm{R}\bm{C}$ with the same eigenvalue $\nu_{i}$, for $i=1,\dots,k$.
Observe that $(\bm{R}\bm{C})^{*}=\bm{C}\bm{R}$, so the non-symmetric matrices $\bm{R}\bm{C}$ and $\bm{C}\bm{R}$ have the same set of eigenvalues, which are also real, as they are the same as the eigenvalues of the symmetric matrix $\bm{R}^{\frac{1}{2}}\bm{C}\bm{R}^{\frac{1}{2}}$. Further, right eigenvectors of $\bm{R}\bm{C}$ are left eigenvectors of $\bm{C}\bm{R}$, and vice versa, left eigenvectors of $\bm{R}\bm{C}$ are right eigenvectors of $\bm{C}\bm{R}$. This finishes the proof. ∎
In particular, if $c_{1}=\dots=c_{k}=c$, then $\nu_{1}=c$. If we further assume the $c_{in}$ versus $c_{out}$ scenario, then in the symmetric case, $\nu_{2}=\dots=\nu_{k}=c\lambda$ with $10<\lambda<$ 1 (see Section 6).
Note that under the condition (14) of Section 3, the eigenvalues of $\bm{R}\bm{C}$ and $\bar{\bm{A}}$ may differ within a factor approaching 1.
The (random) adjacency matrix $\bm{A}$ of (the random graph) $G_{n}$ coming from the $SBM_{k}$ model is $\bm{A}={\bar{\bm{A}}}+\bm{W}$, where $\bm{W}$ is an appropriate (random) error matrix and all the matrices are $n\times n$ symmetric (though we do not use an index $n$ for this fact). We can achieve that the matrix $\bm{A}$ contains 1’s in the $(a,b)$-th block with probability $p_{ab}$, and 0’s otherwise. Indeed, for indices $1\leq a<b\leq k$ and $i\in V_{a}$, $j\in V_{b}$ let
$w_{ji}=w_{ij}:=\left\{\begin{array}[]{ll}1-p_{ab}&\mbox{with probability }\quad p_{ab}\\ -p_{ab}&\mbox{with probability }\quad 1-p_{ab}\end{array}\right.$ (19)
be independent random variables; further, for $a=1,\dots,k$ and $i,j\in V_{a}$ $(i\leq j)$ let
$w_{ji}=w_{ij}:=\left\{\begin{array}[]{ll}1-p_{aa}&\mbox{with probability }\quad p_{aa}\\ -p_{aa}&\mbox{with probability }\quad 1-p_{aa}\end{array}\right.$ (20)
be also independent, otherwise $\bm{W}$ is symmetric. If self-loops are not allowed, then $w_{ii}=-p_{aa}$ with probability 1 could be defined for $i\in V_{a}$. It won’t change significantly the subsequent calculations.
This $\bm{W}$ is not a Wigner noise (see [6]) as it does not have a nested structure. However, its entries are uniformly bounded with zero expectation and variance
$p_{ab}(1-p_{ab})=\frac{c_{ab}(n-c_{ab})}{n^{2}},\quad i\in V_{a},j\in V_{b}.$
We should estimate the spectral norm of $\bm{W}$ with its $L_{1}$-norm so that to apply the Weyl’s perturbation theorem, see [3, 30]. It may help that the expectation of any row-sum of $\bm{W}$ is 0, and the variance of the $i$th row-sum is
$\sum_{b=1}^{k}n_{b}p_{ab}(1-p_{ab})=\sum_{b=1}^{k}r_{b}c_{ab}(1-\frac{c_{ab}}{n})\leq c_{a},\quad a=1,\dots k.$
Therefore, by the Chebyshev’s inequality,
$\mathbb{P}(i\textrm{th row sum}\geq\varepsilon)\leq\frac{c_{a}}{\varepsilon^{2}}.$
Also, if $i\in V_{a}$ and $j\in V_{b}$, then $\mathbb{E}|w_{ij}|=2p_{ab}(1-p_{ab})$, and so, for $i\in V_{a}$:
$\mathbb{E}(\sum_{j=1}^{n}|w_{ij}|)=2\sum_{b=1}^{k}r_{b}c_{ab}(1-\frac{c_{ab}}{n})\leq 2c_{a}.$
Further, if $i\in V_{a}$ and $j\in V_{b}$, then ${\mathrm{Var}}|w_{ij}|=p_{ab}(1-p_{ab})-4p_{ab}^{2}(1-p_{ab})^{2}$ (of course, ${\mathrm{Var}}|w_{ij}|\leq{\mathrm{Var}}(w_{ij})$). Therefore, for $i\in V_{a}$:
${\mathrm{Var}}(\sum_{j=1}^{n}|w_{ij}|)=\sum_{b=1}^{k}r_{b}c_{ab}(1-\frac{c_{ab}}{n})-4\sum_{b=1}^{k}r_{b}c_{ab}\frac{c_{ab}}{n}(1-\frac{c_{ab}}{n})^{2}\leq c_{a}.$
Consequently, if $i\in V_{a}$, then
$\mathbb{P}(|\sum_{j=1}^{n}|w_{ij}|-\mathbb{E}\sum_{j=1}^{n}|w_{ij}||<\varepsilon)\geq 1-\left(\frac{\sqrt{c_{a}}}{\varepsilon}\right)^{2}.$
The right hand side is positive for $\varepsilon\geq\sqrt{c_{a}}$ if $i\in V_{a}$. So, there can be $\max_{a}(2c_{a}+\sqrt{c_{a}})$ fluctuations between the eigenvalues of $\bar{\bm{A}}$ and $\bm{A}$, and large deviation rules do not work here.
Note that this $\bm{W}$ is approximately $\frac{1}{\sqrt{n}}$ times a Wigner noise, and a “semicircle law” is also valid with radius of constant order: $2\sigma=2\max_{a,b}\sqrt{p_{ab}(1-p_{ab})}\leq 1$, see [13]. Now $\bar{\bm{A}}$ is the finite rank ($k$) perturbation on this “deformed Wigner noise”, and if $\lambda_{max}({\bar{\bm{A}}})\sim\lambda_{max}(\bm{B})>1$, then the spectrum of $\bm{A}$ is out of the semicircle.
Summarizing, the eigenvalues of the inflated matrix $\bar{\bm{A}}$ (same as those of $\bm{R}\bm{C}$) are of constant order, but by the Weyl’s perturbation theorem, a constant order of perturbation ($\bm{W}$) can take them farther from those of the true adjacency matrix ($\bm{A}$) with distance $\max_{a}(2c_{a}+\sqrt{c_{a}})$ than from the leading eigenvalues of $\bm{B}$. The same is true for the relation between the $k$ leading eigenvalues of the expected adjacency matrix $\mathbb{E}(\bm{A})$ and $\bm{A}$.
As we saw, $\bm{B}$’s leading eigenvalues are within a factor $\max_{a}\frac{c}{c_{a}}$ from those of $\bm{R}\bm{C}$. In this way, the additive errors can be larger than the multiplicative ones. This is also supported by the theory of deformed Wigner matrices [13].
As for the eigenvectors, in the $SBM_{k}$ model, the eigenvectors of the expected adjacency matrix are $n$-dimensional step-vectors having $k$ different coordinates. If $\mathbf{u}$ denotes such an eigenvector, then the corresponding $2m$-dimensional inflated vector $\bm{End}\,\mathbf{u}$ has the same coordinates with multiplicities equal to the node-degrees $(k\leq n\leq m)$. Those are also step-vectors on $k$ steps, and by Proposition 2, are close to the corresponding eigenvector of $\bm{B}$ (with the close eigenvalue). Therefore, the deflated vector is
${\bm{End}}^{*}{\bm{End}}\,\mathbf{u}=({\bm{End}}\,\mathbf{u})^{in}$
and, by Equation (4), $\bm{D}^{-1}\,({\bm{End}}\,\mathbf{u})^{in}$ approximates $\mathbf{u}$; i.e., the first $n$ coordinates of the corresponding $\bm{K}$-eigenvector, divided with the corresponding node-degrees, will approximate $\mathbf{u}$.
Recall Proposition 2, according to which the leading positive real eigenvalues of $\bm{B}$ (those of $\bm{K}$) are close to those of the expected adjacency matrix. In case of the conditions of Proposition 2, the corresponding normalized eigenvectors of $\bm{B}$ have an inner product with the standardized $\bm{End}$-transforms of the analogous eigenvectors of the expected adjacency matrix very close to 1. But the eigenvectors of the latter one are step-vectors, and the eigenvectors of the former have a large inner product with them, so they are within a “small” distance from the subspace spanned by the two sets of the $k$ leading eigenvectors. This implies that without knowing the model parameters, the $k$-variance of the $k$-dimensional representatives of the edges is “small”. In this way, we obtain $k$ clusters of the bidirected edges, but via the majority of their end-nodes we can classify the majority of the nodes too.
To decrease the computational complexity, we can as well use the right eigenvectors of $\bm{K}$ (only one half of them), more precisely, the vectors $\bm{D}^{-1}\,\mathbf{x}_{i}^{in}$ for $i=1,\dots,k$. For the representatives, based on them, the k-means algorithm is applicable.
More precisely, by [29], let $\mathbf{x}$ be a unit-norm eigenvector of $\bm{B}$, corresponding to the eigenvalue $\mu$ that is close to the eigenvalue $\nu$ of the expected adjacency matrix, with corresponding eigenvector $\mathbf{u}\in\mathbb{R}^{n}$. If our graph is from the $SBM_{k}$ model, then (without knowing its parameters) we know that $\mathbf{u}$ is a step-vector with at most $k$ different coordinates. In [29] it is proved that
$\left\langle\mathbf{x},\frac{{\bm{End}}\,\mathbf{u}}{\|{\bm{End}}\,\mathbf{u}\|}\right\rangle\geq\sqrt{1-\varepsilon}\geq 1-\frac{1}{2}\varepsilon,$
where $\varepsilon$ can be arbitrarily “small” with increasing $n$. Consequently,
$\left\|\mathbf{x}-\frac{{\bm{End}}\,\mathbf{u}}{\|{\bm{End}}\,\mathbf{u}\|}\right\|^{2}\leq 2-2(1-\frac{1}{2}\varepsilon)=\varepsilon$
and, as $\mathbf{x}^{in}=\bm{End}^{*}\mathbf{x}$ holds by (5) and ${\bm{End}}^{*}{\bm{End}}=\bm{D}$ holds by (4),
$\left\|{\bm{End}}^{*}\mathbf{x}-{\bm{End}}^{*}\frac{{\bm{End}}\,\mathbf{u}}{\|{\bm{End}}\,\mathbf{u}\|}\right\|^{2}=\left\|\mathbf{x}^{in}-\bm{D}\frac{\mathbf{u}}{\|{\bm{End}}\,\mathbf{u}\|}\right\|^{2}$
also holds. Consequently,
$\left\|\bm{D}^{-1}\mathbf{x}^{in}-\frac{\mathbf{u}}{\|{\bm{End}}\,\mathbf{u}\|}\right\|^{2}\leq\|\bm{D}^{-1}{\bm{End}}^{*}\|^{2}\varepsilon\leq\varepsilon.$
Indeed, the largest eigenvalue of $(\bm{D}^{-1}{\bm{End}}^{*})({\bm{End}}\,\bm{D}^{-1})=\bm{D}^{-1}\bm{D}\bm{D}^{-1}=\bm{D}^{-1}$ is $\max_{i}\frac{1}{d_{i}}$, so the largest singular value (spectral norm) of $\bm{D}^{-1}{\bm{End}}^{*}$ is estimated from above with $\left(\max_{i}\frac{1}{d_{i}}\right)^{\frac{1}{2}}$. Therefore,
$\|\bm{D}^{-1}{\bm{End}}^{*}\|^{2}\leq\max_{i}\frac{1}{d_{i}}=\frac{1}{\min_{i}d_{i}}\leq 1.$
This results can be summarized in the following theorem.
Theorem 2.
Assume that the expected adjacency matrix of the underlying random graph on $n$ nodes and $m$ edges has rank $k$ with $k$ single eigenvalues and corresponding unit-norm eigenvectors $\mathbf{u}_{1},\dots,\mathbf{u}_{k}\in\mathbb{R}^{n}$. Assume that the non-backtracking matrix $\bm{B}$ of the random graph has $k$ structural eigenvalues (aligned with those of the expected adjacency matrix) with eigenvectors $\mathbf{x}_{1},\dots,\mathbf{x}_{k}\in\mathbb{R}^{2m}$ such that
$\left\langle\mathbf{x}_{j},\frac{{\bm{End}}\,\mathbf{u}_{j}}{\|{\bm{End}}\,\mathbf{u}_{j}\|}\right\rangle\geq\sqrt{1-\varepsilon},\quad j=1,\dots,k.$
Then for the transformed vectors $\bm{D}^{-1}\mathbf{x}_{j}^{in}\in\mathbb{R}^{n}$, the relation
$\sum_{j=1}^{k}\left\|\bm{D}^{-1}\mathbf{x}_{j}^{in}-\frac{\mathbf{u}_{j}}{\|{\bm{End}}\,\mathbf{u}_{j}\|}\right\|^{2}\leq k\varepsilon$ (21)
holds, where $\bm{D}$ is the degree-matrix of the underlying graph.
Corollary 1.
If $\mathbf{u}_{j}$’s are step-vectors on $k$ steps (e.g., if our graph comes from the $SBM_{k}$ model), then the $k$-variance of the node representatives
$(\frac{1}{d_{i}}x_{1i}^{in},\dots,\frac{1}{d_{i}}x_{ki}^{in}),\quad i=1,\dots,n$
is estimated from above with $k\varepsilon$ too. Indeed, the $k$-variance is the minimum squared distance between the subspace spanned by the vectors $\bm{D}^{-1}\mathbf{x}_{j}^{in}$ $(j=1,\dots,k)$ and that of the $k$-step vectors, see [5] for the proof. Since in the $SBM_{k}$ model, $\mathbf{u}_{j}$’s are also step-vectors on the same $k$ steps (as was proved in Proposition 3), this $k$-variance is less than the left-hand side of inequality (21), so it is at most $k\varepsilon$.
Also, if the underlying graph is sparse and the conditions of Theorem 1 of [29] hold, then $\varepsilon>0$ can be arbitrarily small with increasing $n$. To minimize the objective function of the k-means algorithm, there are polynomial time approximating schemes, e.g., [28], so without knowing the model parameters and the vectors $\mathbf{u}_{j}$s, we can find the clusters with the help of the leading $\bm{B}$-eigenvectors only.
It is important that here $k$ is the number of the structural eigenvalues of $\bm{B}$, and the expected adjacency matrix must have a “good” $k$-rank approximation.
This gives rise to a future research on sparse spectral clustering.
Remark 1.
In case of an unweighted graph, the $n$-dimensional vector $\mathbf{x}_{j}^{in}$, obtained from a right eigenvector of $\bm{B}^{*}$, is the second segment of the right eigenvector of the matrix $\bm{K}$, corresponding to the same eigenvalue. Consequently, by the swapping issues of Section 2, the $n$-dimensional vector $\mathbf{x}_{j}^{in}$, obtained from a right eigenvector of $\bm{B}$, is the first segment of the right eigenvector of $\bm{K}$, for $j=1,\dots,k$. So, we have to perform the spectral decomposition of a $2n\times 2n$ matrix only instead of a $2m\times 2m$ one, which fact has further computational benefit (except for trees, $n\leq m$, but usually $n$ is much smaller than $m$).
5 EM algorithm for estimating the parameters of the $SBM_{k}$ model
Given a simple graph $G=(V,\bm{A})$ ($|V|=n$) with adjacency matrix $\bm{A}$ and an integer $k$ $(1<k<n)$, we are looking for the hidden $k$-partition $(V_{1},\dots,V_{k})$ of the nodes such that
  • nodes are independently assigned to cluster $V_{a}$ with probability $r_{a}$, $a=1,\dots,k$; $\sum_{a=1}^{k}r_{a}=1$;
  • given the cluster memberships, nodes of $V_{a}$ and $V_{b}$ are connected independently, with probability
    $\mathbb{P}(i\sim j\,|\,i\in V_{a},j\in V_{b})=p_{ab},\quad 1\leq a,b\leq k.$
The parameters are collected in the probability vector $\mathbf{r}=(r_{1},\dots,r_{k})$ and the $k\times k$ symmetric probability matrix $\bm{P}=(p_{ab})$. We could as well consider the affinity matrix, but it does not make much difference as $n$ is fixed now.
Here the $n\times n$ symmetric, 0-1 adjacency matrix $\bm{A}=(a_{ij})$ of $G$ represents a statistical sample. There are no loops, so the diagonal entries are zeros, whereas the independent off-diagonal entries are considered as an i.i.d. sample. Based on $\bm{A}$, we want to estimate the parameters $\mathbf{r}$ and $\bm{P}$ of the above $SBM_{k}$ model.
Here $\bm{A}$ is the incomplete data specification as the cluster memberships are missing. Therefore, it is straightforward to use the EM algorithm for parameter estimation from incomplete data.
First we complete our data matrix $\bm{A}$ with latent membership vectors $\bm{\Delta}_{1},\dots,\bm{\Delta}_{n}$ of the nodes that are $k$-dimensional i.i.d. $Poly(1,\mathbf{r})$ (multinomially distributed) random vectors. More precisely, $\bm{\Delta}_{i}=(\Delta_{1i},\dots,\Delta_{ki})$, where $\Delta_{ai}=1$ if $i\in V_{a}$ and zero otherwise. Thus, the sum of the coordinates of any $\bm{\Delta}_{i}$ is 1, and $\mathbb{P}(\Delta_{ai}=1)=r_{a}$.
The likelihood is the square-root of
$\prod_{a=1}^{k}\prod_{i=1}^{n}\prod_{b=1}^{k}[p_{ab}^{\sum_{j:\,j\neq i}\Delta_{bj}a_{ij}}\cdot(1-p_{ab})^{\sum_{j:\,j\neq i}\Delta_{bj}(1-a_{ij})}]^{\Delta_{ai}},$ (22)
which is maximized by the following iteration.
Starting with initial parameter values $\mathbf{r}^{(0)}$, $\bm{P}^{(0)}$ and membership vectors $\bm{\Delta}_{1}^{(0)},\dots,\bm{\Delta}_{n}^{(0)}$, the $t$-th step of the iteration is the following ($t=1,2,\dots$).
  • E-step: we calculate the conditional expectation of each $\Delta_{i}$ conditioned on the model parameters and on the other cluster assignments obtained in step $t-1$ and collectively denoted by $M^{(t-1)}$. By the Bayes rule, the responsibility of node $i$ for cluster $a$ is
    $r_{ai}^{(t)}$ $=\mathbb{E}(\Delta_{ai}\,|\,M^{(t-1)})$
    $=\frac{\mathbb{P}(M^{(t-1)}|\Delta_{ai}=1)\cdot r_{a}^{(t-1)}}{\sum_{b=1}^{k}\mathbb{P}(M^{(t-1)}|\Delta_{bi}=1)\cdot r_{b}^{(t-1)}}$
    ($a=1,\dots,k;\,i=1,\dots,n$). For each $i$, $r_{ai}^{(t)}$ is proportional to the numerator, where
    $\mathbb{P}(M^{(t-1)}|\Delta_{ai}=1)$
    $=$ $\prod_{b=1}^{k}(p_{ab}^{(t-1)})^{\sum_{j\neq i}\Delta_{bj}^{(t-1)}a_{ij}}\cdot(1-p_{ab}^{(t-1)})^{\sum_{j\neq i}\Delta_{bj}^{(t-1)}(1-a_{ij})}$
    is the part of the likelihood (22) that affects node $i$ under the condition $\Delta_{ai}=1$.
  • M-step: we maximize the truncated binomial likelihood
    $p_{ab}^{\sum_{i\neq j}r_{ai}^{(t)}r_{bj}^{(t)}a_{ij}}\cdot(1-p_{ab})^{\sum_{i\neq j}r_{ai}^{(t)}r_{bj}^{(t)}(1-a_{ij})}$
    with respect to the parameter $p_{ab}$, for all $a,b$ pairs separately. Obviously, the maximum is attained when the following estimators of $p_{ab}$’s comprise the symmetric matrix $\bm{P}^{(t)}$: $p_{ab}^{(t)}=\frac{\sum_{i,j:\,i\neq j}r_{ai}^{(t)}r_{bj}^{(t)}a_{ij}}{\sum_{i,j:\,i\neq j}r_{ai}^{(t)}r_{bj}^{(t)}}$ $(1\leq a\leq b\leq k)$, where edges connecting nodes of clusters $a$ and $b$ are counted fractionally, multiplied by the membership probabilities of their endpoints.
The maximum likelihood estimator of $\mathbf{r}$ in the $t$-th step is $\mathbf{r}^{(t)}$ of coordinates $r_{a}^{(t)}=\frac{1}{n}\sum_{i=1}^{n}r_{ai}^{(t)}$ ($a=1,\dots,k$), while that of the membership vector $\bm{\Delta}_{i}$ is obtained by discrete maximization: $\Delta_{ai}^{(t)}=1$ if $r_{ai}^{(t)}=\max_{b\in\{1,\dots,k\}}r_{bi}^{(t)}$ and 0, otherwise. (In case of ties, nodes are classified arbitrarily.) This choice of $\mathbf{r}$ will increase (better to say, not decrease) the likelihood function. Note that it is not necessary to assign nodes uniquely to the clusters, the responsibility $r_{ai}$ of a node $i$ can as well be regarded as the intensity of node $i$ belonging to cluster $a$, akin to the fuzzy clustering.
According to the general theory of the EM algorithm [15], in exponential families (as in the present case), convergence to a local maximum can be guaranteed (depending on the starting values), but it runs in polynomial time in the number of nodes $n$. However, the speed and limit of the convergence depend on the starting clustering, which can be chosen by means of preliminary application of some nonparametric multiway cut algorithm.
6 Special cases and $\beta$-percolation
In the most special “symmetric case”, the transition matrix is $\bm{T}=\bm{G}\bm{R}\bm{C}=\frac{1}{ck}\bm{C}$, where $\bm{C}$ contains $c_{in}$ in its main diagonal and $c_{out}$, otherwise; see the BP method of Section 3. $\bm{T}$ is a stochastic matrix, so its largest eigenvalue is 1 with corresponding eigenvector $\mathbf{1}$ (the all 1’s vector). The other eigenvalue is
$\lambda=\frac{c_{in}-c_{out}}{kc}$ (23)
with multiplicity $k-1$. In the assortative case, $\lambda>0$; further, $\lambda<1$, as in the symmetric case
$c=\frac{c_{in}+(k-1)c_{out}}{k}$
holds. Consequently, the eigenvalues of $\bm{R}\bm{C}=c\bm{T}$ are $c$ and $c\lambda$, latter one has multiplicity $k-1$.
With the BP method of [22], that treats this special case, the eigenvalues of $\bm{B}\otimes\bm{T}$ greater than 1 are considered that boils down to the condition $c\lambda^{2}>1$, which gives the Kesten–Stigum threshold (10).
Now assume that $G_{n}\in SBM_{k}^{\beta}$. Then the $k\times k$ probability matrix is $(p_{ab})=\beta\frac{1}{n}\bm{C}=\frac{\beta\bm{C}}{n}$. Consequently, $\bm{C}$ and $c$ is multiplied by $\beta$, but $\bm{T}=\bm{G}\bm{R}\bm{C}$ remains unchanged. So we consider $\beta c\bm{T}$ as for the model side, but the underlying graph and its $\bm{B}$ is the same as before. Therefore, the eigenvalues of $\bm{B}\otimes\beta c\bm{T}=\beta c(\bm{B}\otimes\bm{T})$ should be greater than $c$ if a non-trivial solution is required:
$\beta\,\lambda(\bm{B}\otimes c\bm{T})>c.$
If the eigenvalues of $\bm{B}$ and $c\bm{T}$ are aligned, then this gives that $\lambda(\bm{B})>\frac{\sqrt{c}}{\sqrt{\beta}}$ is needed for detectability; equivalently,
$\beta>\frac{c}{\lambda^{2}(B)}=\left(\frac{\sqrt{c}}{\lambda(\bm{B})}\right)^{2}.$ (24)
This is in accord with the fact, that in the $k=1$ case, in the Erdős–Rényi model, when the largest eigenvalue of $\bm{B}$ is around $c$, then $\beta=\frac{1}{\mu_{1}}$ is the percolation threshold, see [26]. In the multiclass scenario, $\frac{c}{\mu_{i}^{2}}$ are further phase transitions, leading to $i$ clusters, for $i=1,\dots,k_{0}$ until $\mu_{k_{0}}\geq\sqrt{c}$, but $\mu_{k_{0}+1}<\sqrt{c}$. Also note that inequality (24) has relevance only if $\lambda(\bm{B})>\sqrt{c}$, so eigenvalues of $\bm{B}$ greater than $\sqrt{c}$ give the phase transitions.
It is also in accord with the findings in the forthcoming $k=2$ case. Let us consider the symmetric $SBM_{2}$ model: $r_{1}=r_{2}=\frac{1}{2}$, $c_{1}=c_{2}=\frac{c_{in}+c_{out}}{2}=c$. Then condition (10) in the assortative case boils down to
$c_{in}-c_{out}=|c_{in}-c_{out}|>2\sqrt{c}=\sqrt{2}\sqrt{c_{in}+c_{out}}.$ (25)
If $\bm{C}$ and $c$ are multiplied with $\beta$, we get $\beta(c_{in}-c_{out})>\sqrt{2}\sqrt{\beta(c_{in}+c_{out})}$. This means that
$c_{in}-c_{out}>\sqrt{2}\frac{\sqrt{c_{in}+c_{out}}}{\sqrt{\beta}}.$
Since $\beta<1$, the right hand side gives a higher lower threshold than $\beta=1$ in Equation (25). This is also equivalent to
$\beta>\frac{2(c_{in}+c_{out})}{(c_{in}-c_{out})^{2}},$ (26)
which makes sense if $\frac{2(c_{in}+c_{out})}{(c_{in}-c_{out})^{2}}<1$, i.e., if
$c_{in}-c_{out}>2\sqrt{c}=\sqrt{2}\sqrt{c_{in}+c_{out}},$
in accord with (25). So, until equality is attained in (26), additional $\beta$-percolation can give detectable clusters too.
In the special case when $c_{1}=\dots=c_{k}=c$, the above is also equivalent to
$c\lambda>\sqrt{c},$
where $c\lambda$ is the second largest eigenvalue (of multiplicity $k-1$) of $c\bm{T}=\bm{R}\bm{C}$, and it is closely related to the eigenvalues $\mu_{2},\dots,\mu_{k}$ of $\bm{B}$.
The above is also in accord with (24), since the threshold for $\beta$ is
$\beta>\frac{c}{\mu_{2}^{2}}=\frac{c}{(\frac{c_{in}-c_{out}}{2})^{2}}=\frac{\frac{c_{in}+c_{out}}{2}}{\frac{(c_{in}-c_{out})^{2}}{4}}=\frac{2(c_{in}+c_{out})}{(c_{in}-c_{out})^{2}},$
which is the same as inequality (26).
Observe that inequality (10) implies
$\beta|c_{in}-c_{out}|>k\sqrt{\beta c},$
which yields
$\beta>\frac{k^{2}c}{(c_{in}-c_{out})^{2}}=\frac{k[c_{in}+(k-1)c_{out}]}{(c_{in}-c_{out})^{2}}=\frac{c}{(c\lambda)^{2}}=\left(\frac{\sqrt{c}}{\lambda}\right)^{2}$
in the symmetric case, where $\lambda$ is the second largest eigenvalue of $\bm{R}\bm{C}$ with multiplicity $k-1$ and it is aligned with $\mu_{2},\dots,\mu_{k}$. So in the symmetric case, as the second largest eigenvalue of $\bm{R}\bm{C}$ has multiplicity $k-1$, it gives rise to $k-1$ very close phase transitions. In other cases, those can be farther apart.
Another implication is that
$\sqrt{\beta}|c_{in}-c_{out}|>k\sqrt{c},$
where the left-hand side decreases with $\beta<1$. It can be decreased until equality is attained, when $\beta=\frac{c}{\lambda^{2}}$.
In the general scenario, it means that at $\frac{c}{\mu_{i}^{2}}$, for $i=2,\dots,k$, there are phase transitions for $\beta$ producing $2,\dots,k$ detectable clusters:
$|c_{in}-c_{out}|>k\frac{\sqrt{c}}{\sqrt{\beta}}>k\sqrt{c},$
so there is a leverage of the percolation threshold with $1/\sqrt{\beta}$.
In the $c_{in}$ versus $c_{out}$ scenario, $c_{1}=\dots=c_{k}=c$ is equivalent to $r_{1}=\dots=r_{k}$, and so,
$\beta|c_{in}-c_{out}|>k\sqrt{\beta c},\quad|c_{in}-c_{out}|>k\frac{\sqrt{c}}{\sqrt{\beta}}$
which allows more detectable clusters if $\beta$ is increased, see the subsequent illustrations.
Eventually, this phenomenon is illustrated by simulations. First a random graph was generated on $n=900$ nodes, with parameter matrices $\bm{R}=\mathrm{diag}(\frac{3}{10},\frac{1}{3},\frac{11}{30})$ and
$\bm{C}=\begin{pmatrix}30&12&10\\ 12&32&9\\ 10&9&27\end{pmatrix}.$
Then we percolated (retained) the edges with probability $\beta$. For small $\beta$’s the graph was not connected, which fact caused the eigenvalue $1$ of the matrix $\bm{K}$ to have several geometric, and even more algebraic multiplicities. With increasing $\beta$, Figures 123, and 4 show the adjacency matrices and the $\bm{K}$-spectrum of the $\beta$-percolated random graph, when 1,2, and 3 structural (positive real) eigenvalues emerge out of the bulk (within the encircled region). This happens about when $\beta$ exceeds $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$, but the phase transitions are not strict, as the alignment of the structural eigenvalues of the non-backtracking matrix $\bm{B}$ (those of $\bm{K}$) and those of the model matrix $\bm{R}\bm{C}$ is just asymptotic (for “large” $n$) and guaranteed only when the expected degrees of the clusters are approximately the same.
(a) Adjacency matrix (black: 1, white: 0).
(b) Spectrum of matrix $\bm{K}$.
Figure 1: $\beta_{1}<\beta=0.191<\beta_{2}$
(a) Adjacency matrix (black: 1, white: 0).
(b) Spectrum of matrix $\bm{K}$.
Figure 2: $\beta=\beta_{3}=0.505$
(a) Adjacency matrix (black: 1, white: 0).
(b) Spectrum of matrix $\bm{K}$.
Figure 3: $\beta_{3}<\beta=0.7$
(a) Adjacency matrix (black: 1, white: 0).
(b) Spectrum of matrix $\bm{K}$.
Figure 4: $\beta=1$
Next a random graph was generated on $n=900$ nodes, with parameter matrices $\bm{R}=\mathrm{diag}(\frac{35}{107},\frac{42}{107},\frac{30}{107})$ and
$\bm{C}=\begin{pmatrix}30&11.28&7.728\\ 11.28&25&10.36\\ 7.728&10.36&35\end{pmatrix}$
constructed so that the average degrees of the clusters be the same, i.e., $c_{a}=\sum_{b=1}^{3}r_{b}c_{ab}$ is the same for $a=1,2,3$.
Figures 567, and 8 show the adjacency matrices and the $\bm{K}$-spectrum of the $\beta$-percolated random graph, when 1,2, and 3 structural (positive real) eigenvalues appear. Here the phase transitions are closer to $\beta_{i}$’s than in the previous generic case.
(a) Adjacency matrix (black: 1, white: 0).
(b) Spectrum of matrix $\bm{K}$.
Figure 5: $\beta_{1}<\beta=0.183<\beta_{2}$
(a) Adjacency matrix (black: 1, white: 0).
(b) Spectrum of matrix $\bm{K}$.
Figure 6: $\beta=\beta_{2}=0.305$
(a) Adjacency matrix (black: 1, white: 0).
(b) Spectrum of matrix $\bm{K}$.
Figure 7: $\beta_{2}<\beta=0.563<\beta_{3}$
(a) Adjacency matrix (black: 1, white: 0).
(b) Spectrum of matrix $\bm{K}$.
Figure 8: $\beta=1$
Summarizing, at $\beta_{1}$, the giant component is formed. In [27], the effect of different seeding patterns is investigated in context of the COVID19 pandemic in Hungary. According to this, below a certain $\beta$, the central seeding is more dangerous than the uniform one. Indeed, below $\beta_{1}$ there are many small components, among which there is a central one. Above that, the uniform seeding is more dangerous than the central one, as it is able to infect the giant component and other components too. However, if the seeds arrive in time, then actions can be done to separate the most infected clusters as follows.
At $\beta_{2}$, there are 2 well-distinguishable clusters, …, at $\beta_{k}$, there are $k$ well-distinguishable ones, since the within-cluster bonds are strengthened. It does not mean that there are no between-cluster bonds, but those are negligible compared to the within-cluster ones. Also, by deleting the between-cluster bonds, the infection can be confined to the clusters. If the seed is in one cluster, the epidemic can be confined to that cluster. So with increasing $\beta$, there is a chance to localize the epidemic to one large city and to save the others.
For example, if $\beta$ passes over $\beta_{2}$, then 2 clusters can be separated, and if the seed is in one of them, then the infection can be localized to that cluster and we can save the other by deleting the connecting bonds (even if we do not delete all, the epidemic will spread to the other cluster with small probability only). If $\beta$ passes over $\beta_{3}$, then 3 clusters can be separated, and if the seed is in one of them, then the infection can be localized to that cluster and we can save the others by deleting the connecting bonds, etc. Above $\beta_{k}$ the situation is similar to that of $\beta=1$, the uniform seeding is dangerous, but seeds can be localized to the first infected cluster. In the symmetric case, $\beta_{2}=\dots=\beta_{k}$, so all these phase transitions occur at the same time from the giant cluster to the $k$-cluster scenario. Multiple transitions are spectacular if the eigenvalues of $\bm{B}$ greater than $\sqrt{c}$ are separated from each other and from $\sqrt{c}$. Therefore, $\beta$ can as well be considered as a tuning parameter. Also note that in the $SBM_{k}$ model when $c_{a}=c$ $(a=1,\dots,k)$, then $\mu_{1}$ is close to $c$ and so, $\beta_{1}$ is close to $\frac{1}{\mu_{1}}$ (w.h.p.). This is always the case if $k=1$, and so, $\frac{c}{\mu_{1}^{2}}=\frac{1}{\mu_{1}}$ as in [26].
7 Conclusions and further perspectives
The sparse stochastic block model behind an expanding graph sequence was considered. Here the spectral properties of the adjacency matrix give less information for the clusters than in the dense case of [9, 20]; instead, the non-backtracking spectra are used. Bond-percolation is considered in two senses: the first is that the within- and between-cluster edge probabilities decrease with $n$, and the second is that edges coming into existence in this way are retained with probability $\beta$. Similarities and differences between the BP and EM algorithms are discussed. Both are based on an iteration algorithm, but former one solves a system of equations with the transmission probabilities between the node pairs ($2mk$ non-linear equations and unknowns, where $k$ is the number of clusters), while keeping the model parameters fixed; latter one, for fixed $k$, estimates the model parameters and the missing memberships of the nodes by using the classical EM iteration for mixtures. Furthermore, via inflation–deflation techniques and results of [29], we are able to classify the nodes of the sparse graph with representation based k-means clustering, which gives rise to a sparse spectral clustering procedure. Simulation results for the so-called $\beta$-percolation are also considered, and strategies are suggested for possible localization of a pandemic.
If the $c_{in}$ versus $c_{out}$ scenario does not hold, then in generic assortative networks there are $k$ different within-cluster and ($k\choose 2$ different between-cluster affinities; former ones, denoted by $c_{w}$’s, are significantly “larger” than the latter ones, denoted by $c_{b}$’s. In this case, $c_{1}=\dots=c_{k}=c$ is not necessarily equivalent to $r_{1}=\dots=r_{k}$, and the Kesten–Stigum threshold should be developed by giving lower bounds for every $c_{w}-c_{b}$ difference, which are positive in assortative networks. Finding such dense diagonal blocks is an important problem of quantum chemistry, for example. In the estimation, discrepancy techniques, see [7, 8] may help.
Author contributions
The theoretical parts of the paper with theorems and proofs were written by Marianna Bolla. The Python code for calculating the non-backtracking spectra and generating random graphs from the $SBM_{k}$ model was written by Daniel Zhou. He also performed the $\beta$-percolation on the so generated random graphs.
Acknowledgments
This research was partly done under the auspices of the Budapest Semesters in Mathematics program, in the framework of a research course in the spring semester 2023, with the supervision of the first author and participation of the second author and Josephine Walk who helped to organize the early simulation results of the second author into digestible form. We are indebted to László Lovász for valuable conversations on this topic. We also thank Tamás Móri, Katalin Friedl and Daniel Keliger for their useful remarks.